OpenStudy (anonymous):

I Am Stuck PLEASE! HELP! Q. What is 9/18 reduced to it's simplest form?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

1/2

OpenStudy (sleepyjess):

Don't just give the answer please

OpenStudy (anonymous):

9 is one half of 18 to so it would be 1/2 just like 1 is one half of 2, which is lowest and simplest form

OpenStudy (anonymous):

divide the numerator and denominator by g.c.d ,in this case by 9

OpenStudy (studygurl14):

\[\frac{ 9 }{ 18 } \div 3 = \frac{ 3 }{ 6 } \] \[\frac{ 3 }{ 6 } \div 3 = \frac{ 1 }{ 2 }\] Or, you can go straight from 9/18 to 1/2... \[\frac{ 9 }{ 18 } \div 9 = \frac{ 1 }{ 2 }\]

OpenStudy (studygurl14):

what @surjithayer was good

OpenStudy (anonymous):

Thanks HatchetLover StudyGirl14 and Sirjithayer and Jmanjeter2! Y'all All Helped Me!